40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 4 02:13:02 2020, gautam, Update, LSC, Locking updates modifiedOutMat.pngOLTFs.pdf
    Reply  Wed Feb 5 16:35:12 2020, gautam, Update, LSC, Diagnosis plan ALSnoise.pdf
       Reply  Thu Feb 6 01:16:58 2020, gautam, Update, LSC, Diagnosis results IMCspec_ALS.pdfIMCspec_ALS_smallSpan.pdf
    Reply  Thu Feb 6 01:25:50 2020, gautam, Update, LSC, Locking updates semiStableArms.pngarmAngStability.png
       Reply  Fri Feb 7 02:24:24 2020, gautam, Update, LSC, Some short notes 
Message ID: 15185     Entry time: Tue Feb 4 02:13:02 2020     Reply to this: 15188   15192
Author: gautam 
Type: Update 
Category: LSC 
Subject: Locking updates 

Summary:

The CARM-->RF transition remains out of reach. No systematic diagnosis scheme comes to mind.

Details:

  • Config is PRFPMI, SRM is misaligned macroscopically.
  • PRMI can easily be locked with 3f signals while CARM is offset from resonance. Aided by DAFI, I turned on the PR violin filter in the BS output section to prevent it from ringing up, making the lock much more robust.
  • When the CARM offset is reduced
    • POP22 level dips and sometimes goes negative - i don't see this in my simple simulations. POP22 is the trigger signal for MICH/PRCL loops, so to prevent the PRMI lockloss, I mix in some POPDC into the trigger matrix element.
    • Once the circulating power exceeds ~10, the ALS noise apparently increases.
    • The arms "buzz" through resonance, but the power fluctuation is nearly 0-200 in TRX/TRY, corresponding to several CARM linewidths, but all the out-of-loop ALS noise measurements have me believe that we are close to the CARM linewidth in noise. So we should only see ~factor of 2 fluctuation in power.
    • The RF error signal for CARM (=REFL 11) doesn't show any features that i can use to aid the transition / diagnose what is going on systematically.
  • Koji suggested changing the actuation for CARM from MC2 to the ETMs, and check if the MC OSEMs witness the excess motion at small CARM offsets
    • The ALS transition is scripted, so I had to make a modified version that accommodates this changed actuation scheme.
    • The usual CARM-->MC2 matrix element is -1. 
    • The frequency actuation strength of MC2 is ~3x that of the ETMs. Additionally, ETMX has 5x the series resistance of ETMY. So I used the output matrix elements shown in Attachment #1 so as to get the same loop UGF with the same loop gains elsewhere in the chain. Confirmed the actuation strength is the same using the sensing matrix infrastructure and comparing line heights.
    • Attachment #2 shows the measured UGF - both CARM and DARM look okay to me.
  • With this new ALS output matrix actuation scheme, I was able to make it to PRMI + arms on zero offset a couple of times tonight, but the drifting input alignment makes the PRMI lock not so robust anymore.

TBC. Mercifully at least the shaker stayed still tonight.

Attachment 1: modifiedOutMat.png  10 kB  | Hide | Hide all
modifiedOutMat.png
Attachment 2: OLTFs.pdf  56 kB  Uploaded Thu Feb 6 16:06:27 2020  | Hide | Hide all
OLTFs.pdf
ELOG V3.1.3-